OpenStudy (anonymous):

7(x − 1) = 5(x + 2) + 1

OpenStudy (precal):

distribute 7 and then distribute 5

OpenStudy (anonymous):

Solve: 7(x − 1) = 5(x + 2) + 1 9 18 2 4

OpenStudy (precal):

I am helping you by telling you how to do the problem part of the code of conduct here at OS

OpenStudy (anonymous):

open the brackets 7x - 7 = 5x +10 +1 2x = 18 x = 9